22west:

Which number should each side of the equation 34x=9 be multiplied by to produce the equivalent equation of x = 12? −34 14 43 4

demolisher1224:

Hey i got a trick @ThisGirlPretty @dude @Elsa213

22west:

srry the 34 is supposed to be \[\frac{ 3 }{ 4 }\]

dude:

Oh, I've been summoned Are the options fractions?

22west:

yes the -34 is \[-\frac{ 3 }{ 4 }\]

dude:

Just that one?

22west:

and the 14 and 43 are fractions

dude:

Ah okay

22west:

here is a snip

22west:

1 attachment
22west:

there

22west:

HI ELSA

Elsa213:

Hai there~ e.e

dude:

In order to get rid of fractions you must multiply by its inverse

22west:

huh?

dude:

Do you know what the inverse of a fraction is?

22west:

negatori

22west:

no i dont

dude:

Flip the fraction :P \(\frac{1}{2}=\frac{2}{1}\)

22west:

pshhh i knew that

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Flip the fraction :P \(\frac{1}{2}=\frac{2}{1}\) \(\color{#0cbb34}{\text{End of Quote}}\) They are NOT equal, that is my mistake \(\frac{1}{2}≠\frac{2}{1}\)

22west:

Flip the fraction :P

dude:

Right, so what is the inverse of \(\frac34\)?

22west:

\[\frac{ 4 }{ 3 }\]

dude:

That is your answer

22west:

oh otay fo mo

22west:

four more

22west:

Solve for c. −5/8c=20 What is the answer? −34 −32 26 28

22west:

bra ya gotta learn slang bra

dude:

Sure, anyway Any ideas on what to do first?

22west:

idea

22west:

first eat a buritto

dude:

22west, you you want help?

22west:

ya i want halp

22west:

bidkwtdf

22west:

but i dont know what to do first

22west:

i FURgot

dude:

Just like in the last problem, we can multiply by its inverse \(-\boxed{\frac58}c=20\)

22west:

ok

dude:

Just like in the last problem, we can multiply by its inverse \(-\boxed{\frac58}c=20\) We can multiply by \(\frac85\) on both sides

dude:

Or we can make this more simple and multiply by 8 on both sides and then divide by 5 on both sides

22west:

umm

22west:

plz tell me the answer

22west:

everyone else does it

dude:

Lets go simple \[-\frac{5}{8}\color{red}{\cdot8}~c=20\color{red}{\cdot 8}\]

dude:

That goes against the rules here

dude:

@Ultrilliam

22west:

he does not need to be here plz

dude:

I can only assist you throughout the problem. I cannot directly give you the answer

22west:

everyone does it ultri

Ultrilliam:

1516830232-5a68f6d681b9a1156f78e000-image.png

22west:

i always read that stuff and i did not see it

22west:

ok i never do